# Computing Floquet quasienergies in finite and extended systems: Role of electromagnetic and quantum-geometric gauges

### v1

@@ III.2.2 Results
Figure 1: Quasienergy spectrum of the first 20 states in the first quasienergy Brillouin zone of a system composed by a particle of charge $q$ and mass $m$ in an infinite potential well of width $a$ subjected to a driving given by the electrical field $E(t)=E_{0}\cos(\omega t)$. The frequency $\omega$ is fixed such that $\hbar\omega=0.95\times 3\hbar^{2}\pi^{2}/(2ma^{2})$ and $E_{0}$ is varied in the $[0,-10\hbar\omega/qa]$ range. The red line shows the spectrum calculated using the length gauge matrix elements Eq. (32), $\varepsilon_{n}^{E}$. The blue and black lines show the spectrum calculated using the velocity gauge matrix elements in the form of Eqs. (34) and (35), $\varepsilon_{n}^{V}$ and $\varepsilon_{n}^{V_{2}}$, respectively.
The figure shows three different calculations corresponding to ${H}^{E}_{nm}(t)$, ${H}^{V}_{nm}(t)$ and ${H}^{V_{2}}_{nm}(t)$. The quasi-energy dispersion for ${H}^{E}_{nm}(t)$ and ${H}^{V}_{nm}(t)$ agree up to numerical precision in all the considered driving-field range and for all eigenstates. This offers the numerical verification we were looking for, namely the gauge-invariant nature of quasienergies as a result of the gauge-covariance of the Floquet Hamiltonian derived in Eq. (21). In addition, our results agree completely with the spectrum reported in Fig. 3 of Ref. [34] up to an irrelevant rigid shift of the whole set of quasienergies. In particular, Fig. 1 correctly describes the strong dispersion of the two quasienergies located at $\varepsilon/\hbar\omega\simeq 0.35$ for low fields, which correspond to the lowest two eigenstates that are close to resonance in Eq. (37[34].
On the other hand, the quasienergy spectrum corresponding to $\hat{H}^{V_{2}}_{nm}(t)$ in Fig. 1 completely deviates from the other two curves and features a nonphysical behavior; only in the weak field limit $E_{0}\rightarrow 0$ does it approach the correct value. This offers a clear numerical evidence of the inadequacy of the velocity-gauge expression in Eq. (34) for performing calculations using a finite basis. We have verified that this deviation decreases as the number of eigenstates in the calculation increases, as expected. In App. B we study an improved version of the $\hat{H}^{V_{2}}(t)$ Hamiltonian obtained by adding an expansion that is introduced in the coming section.
@@ IV Extended systems
The particle in a box described in the previous section constitutes an example of a finite system; atoms and molecules are real-life examples of this type. In contrast, electrons in a crystal or atoms trapped in a periodic optical lattice constitute an extended system. In these cases, the potential entering the ground-state Hamiltonian, Eq. (1) generally obeys the periodic property
@@
where $\bm{R}$ is a Bravais lattice vector. When subjected to a time-dependent perturbation, the calculation of the Hamiltonian in this type of system requires extra care. The fundamental reason for the additional difficulty lies in the position operator, whose representation becomes subtle in an extended system [35]. In this section we describe the scheme we adopted to deal with $\hat{\bm{r}}$ in practical Floquet calculations in extended systems.
@@ IV.1 Quantum-geometric gauge and representation of $\hat{\bm{r}}$
This basis transforms as irreducible representations of the Bravais lattice translation operators $\hat{\mathcal{T}}(\bm{R})$, and its labelling corresponds to a vector $\bm{k}$ within the first Brillouin zone (BZ) of the system [44]. In this extended basis, the closure and orthogonality properties read
@@
Let us now come to the position operator. One can show (see App. D) that, in the Bloch representation, the operator can be expressed as [35]
@@
In the language of differential geometry [45], we are analysing a fibre bundle whose base manifold is the $d$-torus (the BZ), the typical fibre is the $N$-dimensional Hilbert space where the |$\ket{u_{n}(\bm{k})}$ are defined and the gauge group is $U(N)$. For each variable $k_{j}$ parametrizing the base manifold, there exists an associated covariant derivative $(i\hat{\mathcal{D}}^{j})$ and connection form $\hat{\xi}^{j}(\bm{k})$. These objects are required to provide a notion of parallel transportation and are solely defined though their respective transformation properties.
@@ IV.2 Perturbative expansion in the velocity gauge
The above considerations need to be taken into account for calculating Floquet quasienergies in extended systems through the Hamiltonian $\hat{H}(t)$ in either form, length gauge [Eq. (32)] or velocity gauge [Eq. (34)]. The most challenging aspect is numerical implementation: taking the derivative of the delta function in Eq. (46) is problematic since it involves the nonlocal difference $\bm{k}-\bm{q}$. It is worth noting that the difficulty faced here is more severe than that encountered in the perturbative analysis of optical responses of solids. In the latter case and taking as reference the length gauge, $\hat{\bm{r}}$ always enters the response coefficients via commutators with an operator $\hat{O}$, which is the well-defined covariant derivative of operator $\hat{O}$ [49, 48, 40]
@@
In the case that concerns us, on the other hand, calculating $\hat{H}(t)$ via the length-gauge expression of Eq. (32) requires computing the position matrix element plainly. Previous works [50, 51, 52, 34] have proceeded by considering just the interband piece of the Berry connection, namely by approximating
@@
While this has the obvious advantage of not having to deal with the delta-function piece in Eq. (46), it appears to have several drawbacks. In first place, information of the intraband contribution to the Berry connection (proportional to $\delta_{nn}$) is completely lost. As for the discarding of the delta-function piece, it seems a rather uncontrolled procedure considering the highly singular nature of this contribution. Finally on a more fundamental level, Eq. (48) maintains gauge-covariance only under a diagonal transformation of the form ${\mathcal{U}}_{nm}(\bm{k})=\delta_{nm}e^{i\phi_{n}(\bm{k})}$ [gauge group $U(1)$], but it is not gauge-covariant under a general transformation $\hat{\mathcal{U}}$ [gauge group $U(N)$].
@@
which is a power series of objects involving the successive application of the quantum-geometric covariant derivative of Eq. (47) on $\hat{H}_{0}$. In a practical calculation one needs to consider a finite number $M$ of expansion terms. Their explicit element-wise expression can be straightforwardly obtained using Eq. (47). In this work we evaluate the above terms by employing the so-called Wannier interpolation technique; a description of the gauge choice and further technical details are given in App. E.
@@ IV.2.1 Quantifying truncation errors
In Fig. 2 we illustrate the error estimates up to $M=4$. Let us first focus on $\lambda<1$, where central dynamical effects such as multiphoton resonances are expected to take place [34]. In this range, the $M=2$ approximation already yields an error below $\sim 10\%$, while for $M>2$ the error goes below $\sim 1\%$. Therefore, it appears reasonable to expect that keeping the first few terms in the expansion of Eq. (49) suffices to provide a fair estimate of this important regime. For $\lambda>1$, on the other hand, the error in Fig. 2 quickly increases above $\sim 10\%$ even for $M=4$. This regime corresponds to the large electric-field or slow driving-frequency limit.
@@ V Graphitic BC2N
In this section we present the results of the calculations of Floquet quasienergies using the extended-system approach of Sec. IV for the case of the material BC2N.
@@ V.1 Technical details
We have conducted an ab-initio analysis for this material. The first step consisted of a density functional theory calculation using the Quantum Espresso [55] package. We employed a $10\times 10\times 10$ k-mesh both for the self-consistent calculation of the ground state as well as for the subsequent non-self-consistent calculation of the energy eigenvalues and Bloch eigenfunctions. The core-valence interaction was treated by means of scalar relativistic projector-augmented-wave pseudopotentials that had been generated with the Perdew-Burke-Ernzerhof exchange-correlation functional. The pseudopotentials were taken from PS library, generated using the atomic code by A. Dal Corso v.6.3, and the energy cutoff for the plane-wave basis expansion was set at 70 Ry. In the next step, we constructed maximally localized Wannier functions using the Wannier90 package [56]. After discarding the 2 lowest-lying bands, we considered a set of 8 disentangled Wannier functions around the Fermi level to span the “inner” energy range [57] $[-3.33,3.35]\;\text{eV}$. As initial projections, we considered $p_{z}$ atom-centered orbitals for all atoms.
In order to calculate the quasienergy spectrum $\varepsilon_{n}(\bm{k})$, we implement the time-discretization method detailed in App. C, with matrix representation of the Hamiltonian $\hat{H}(t_{j})$ given by the different methods discussed in Sec. IV.2: the length gauge, Eq. (48), and the velocity gauge Eq. (49) keeping terms up to $M=1$, $M=2$ and $M=3$. The implementation of the Hamiltonian and all required quantities has been carried out using the SsTC [58] and WannInt [59] packages, which are specially tailored Fortran libraries to perform Wannier post-processing calculations with special emphasis on computing quantities dependent on a large number of external variables (such as the quasienergies, which depend functionally on $\omega$, and $\bm{E}(t)$).
@@ V.2 Band structure
Figure 3: Band structure of mirror-symmetric BC2N along the $k_{z}=0$ $\Gamma-\text{X}-\text{S}-\text{Y}-\Gamma$ path in the BZ and in the $[-9.0,11.5]\;\text{eV}$ energy range centred around the Fermi energy. The displayed bands have a definite parity with respect to the $\mathcal{M}_{x}$ point group operation, $A_{n}$, which is displayed as a coloured over line. The red (blue) overline represents the $+1$ ($-1$) eigenvalue.
@@ V.3.1 General properties and quality of approximations
Figure. 4 shows the quasienergy spectrum of the 8 bands considered in Fig. 3 as a function of the electric field amplitude. The calculations were performed at k-point $\bm{k}=(1/2,1/4,0)$ (crystal coordinates) lying along the X-S line, and fixed frequency $\hbar\omega=0.5\;\text{eV}$, which corresponds to an almost resonant excitation between the highest occupied molecular orbital (HOMO) and the lowest unoccupied molecular orbital (LUMO). The figure includes three results calculated in the velocity gauge obtained by keeping a different number of expansion terms in Eq. (49); up to $M=1$, $M=2$, and $M=3$. In addition, we have also included results calculated using the length-gauge approximation of Eq. (48), employed in other works.
Figure 4 distinguishes three regimes as a function of $E$, which we have separated by vertical lines for the case of the velocity-gauge results as a guide to the eye. The first is the noninteracting regime of small $E$, where quasienergy bands are virtually flat. The second is an intermediate electric-field regime where the onset of dispersion takes place and is the physically most interesting one. Among other features, this regime shows wide, pronounced avoided crossings that mark multiphoton resonances, which we study in detail in Sec. V.3.2 below. The third regime is the continuum limit of large $E$, which corresponds to $\lambda\gg 1$ [c.f. Eq. (50)] wherein transitions between virtually all energy levels are possible, yielding a featureless continuum spectrum.
@@
V.3.2 Avoided crossings, multiphoton resonances and generalized parity
@@
where $A_{n}$ is the parity with respect the $\mathcal{M}_{x}$ operation, as introduced in the discussion of the band structure in Fig. 3. In the case of the HOMO and LUMO states we have $A_{\text{HOMO}}=-A_{\text{LUMO}}=-1$, while the representatives are $m=14$ for the HOMO and $m^{\prime}=15$ for the LUMO state (notice that translation between quasienergy first Brillouin zone is accompanied by a shift in the representative in Fig. 4), hence $P_{\text{HOMO},14}=P_{\text{LUMO},15}$. From the group theory perspective, the quasienergies corresponding to these states cannot cross and therefore undergo an avoided crossing that marks a one-photon resonance since $|m-m^{\prime}|=1$.
@@ VI Summary and Outlook
In this work, we have studied two central quantities of the Floquet theory, namely the effective Floquet Hamiltonian and its eigenvalues, the quasienergies. In the first part of the work, we have analyzed the implications of the electromagnetic gauge freedom in the transformation properties of these quantities. This subject has attracted relatively little attention in the literature so far, and it appears to have been only studied using perturbation theory up to second order in the external field [61, 62]. Here we have explicitly proved the gauge covariance of the effective Floquet Hamiltonian, which ensures the gauge invariance of the quasienergy spectrum to all orders. We have provided a numerical example of the latter by discussing the results obtained in the velocity and length gauges for a particle trapped in a one-dimensional box subjected to a periodic drive.
@@ VII Acknowledgments
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the European Research Council (ERC) grant agreement No 946629, Grants No. PID2021-129035NB-I00 and PID2021- 126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, and from the Basque Government through Grant No. IT1470-22.

### v2

@@ III.2.2 Results
Figure 1: Quasienergy spectrum of the first 20 states in the first quasienergy Brillouin zone of a system composed by a particle of charge $q$ and mass $m$ in an infinite potential well of width $a$ subjected to a driving given by the electrical field $E(t)=E_{0}\cos(\omega t)$. The frequency $\omega$ is fixed such that $\hbar\omega=0.95\times 3\hbar^{2}\pi^{2}/(2ma^{2})$ and $E_{0}$ is varied in the $[0,-10\hbar\omega/qa]$ range. The red line shows the spectrum calculated using the length gauge matrix elements in Eq. (32), $\varepsilon_{n}^{E}$. The blue and black lines show the spectrum calculated using the velocity gauge matrix elements in the form of Eqs. (34) and (35), $\varepsilon_{n}^{V}$ and $\varepsilon_{n}^{V_{2}}$, respectively.
The figure shows three different calculations corresponding to ${H}^{E}_{nm}(t)$, ${H}^{V}_{nm}(t)$, and ${H}^{V_{2}}_{nm}(t)$. The quasi-energy dispersion for ${H}^{E}_{nm}(t)$ and ${H}^{V}_{nm}(t)$ agree up to numerical precision in all the considered driving field ranges and for all eigenstates. This offers the numerical verification we were looking for, namely the gauge-invariant nature of quasienergies as a result of the gauge-covariance of the Floquet Hamiltonian derived in Eq. (21). In addition, our results agree completely with the spectrum reported in Fig. 3 of Ref. [34] up to an irrelevant rigid shift of the whole set of quasienergies. In particular, Fig. 1 correctly describes the strong dispersion of the two quasienergies located at $\varepsilon/\hbar\omega\simeq 0.35$ for low fields, which correspond to the lowest two eigenstates that are close to resonance in Eq. (37[34].
On the other hand, the quasienergy spectrum corresponding to $\hat{H}^{V_{2}}_{nm}(t)$ in Fig. 1 completely deviates from the other two curves and features a nonphysical behavior; only in the weak field limit $E_{0}\rightarrow 0$ does it approach the correct value. This offers a clear numerical evidence of the inadequacy of the velocity gauge expression in Eq. (34) for performing calculations using a finite basis. We have verified that this deviation decreases as the number of eigenstates in the calculation increases, as expected. In Appendix B we study an improved version of the $\hat{H}^{V_{2}}(t)$ Hamiltonian obtained by adding an expansion that is introduced in the coming section.
@@ IV Extended systems
The particle in a box described in the previous section constitutes an example of a finite system; atoms and molecules are real-life examples of this type. In contrast, electrons in a crystal or atoms trapped in a periodic optical lattice constitute an extended system. In these cases, the potential entering the ground state Hamiltonian, Eq. (1) generally obeys the periodic property
@@
where $\bm{R}$ is a Bravais lattice vector. When subjected to a time dependent perturbation, the calculation of the Hamiltonian in this type of system requires extra care. The fundamental reason for the additional difficulty lies in the position operator, whose representation becomes subtle in an extended system [35]. In this section we describe the scheme we adopted to deal with $\hat{\bm{r}}$ in practical Floquet calculations in extended systems.
@@ IV.1 Quantum-geometric gauge and representation of $\hat{\bm{r}}$
This basis transforms as irreducible representations of the Bravais lattice translation operators $\hat{\mathcal{T}}(\bm{R})$, and its labeling corresponds to a vector $\bm{k}$ within the first Brillouin zone (BZ) of the system [44]. In this extended basis, the closure and orthogonality properties read
@@
Let us now come to the position operator. One can show (see Appendix D) that, in the Bloch representation, the operator can be expressed as [35]
@@
In the language of differential geometry [45], we are analyzing a fiber bundle whose base manifold is the $d$-torus (the BZ), the typical fiber is the $N$-dimensional Hilbert space where the |$\ket{u_{n}(\bm{k})}$ are defined and the gauge group is $U(N)$. For each variable $k_{j}$ parametrizing the base manifold, there exists an associated covariant derivative $(i\hat{\mathcal{D}}^{j})$ and connection form $\hat{\xi}^{j}(\bm{k})$. These objects are required to provide a notion of parallel transportation and are solely defined though their respective transformation properties.
@@ IV.2 Perturbative expansion in the velocity gauge
The above considerations need to be taken into account for calculating Floquet quasienergies in extended systems through the Hamiltonian $\hat{H}(t)$ in either form, length gauge [Eq. (32)], or velocity gauge [Eq. (34)]. The most challenging aspect is numerical implementation: taking the derivative of the delta function in Eq. (46) is problematic since it involves the nonlocal difference $\bm{k}-\bm{q}$. It is worth noting that the difficulty faced here is more severe than that encountered in the perturbative analysis of optical responses of solids. In the latter case and taking as reference the length gauge, $\hat{\bm{r}}$ always enters the response coefficients via commutators with an operator $\hat{O}$, which is the well-defined covariant derivative of operator $\hat{O}$ [49, 48, 40]
@@
In the case that concerns us, on the other hand, calculating $\hat{H}(t)$ via the length gauge expression of Eq. (32) requires computing the position matrix element plainly. Previous works [50, 51, 52, 34] have proceeded by considering just the interband piece of the Berry connection, namely by approximating
@@
While this has the obvious advantage of not having to deal with the Dirac delta piece in Eq. (46), it appears to have several drawbacks. In the first place, information of the intraband contribution to the Berry connection (proportional to $\delta_{nn}$) is completely lost. As for the discarding of the Dirac delta piece, it seems a rather uncontrolled procedure considering the highly singular nature of this contribution. Finally on a more fundamental level, Eq. (48) maintains gauge-covariance only under a diagonal transformation of the form ${\mathcal{U}}_{nm}(\bm{k})=\delta_{nm}e^{i\phi_{n}(\bm{k})}$ [gauge group $U(1)$], but it is not gauge-covariant under a general transformation $\hat{\mathcal{U}}$ [gauge group $U(N)$].
@@
which is a power series of objects involving the successive application of the quantum-geometric covariant derivative of Eq. (47) on $\hat{H}_{0}$. In a practical calculation one needs to consider a finite number $M$ of expansion terms. Their explicit element wise expression can be straightforwardly obtained using Eq. (47). In this work we evaluate the above terms by employing the so-called Wannier interpolation technique; a description of the gauge choice and further technical details are given in Appendix E.
@@ IV.2.1 Quantifying truncation errors
In Fig. 2 we illustrate the error estimates up to $M=4$. Let us first focus on $\lambda<1$, where central dynamical effects such as multiphoton resonances are expected to take place [34]. In this range, the $M=2$ approximation already yields an error below $\sim 10\%$, while for $M>2$ the error goes below $\sim 1\%$. Therefore, it appears reasonable to expect that keeping the first few terms in the expansion of Eq. (49) suffices to provide a fair estimate of this important regime. For $\lambda>1$, on the other hand, the error in Fig. 2 quickly increases above $\sim 10\%$ even for $M=4$. This regime corresponds to the large electric field or slow driving frequency limit.
@@ V Graphitic BC2N
In this section we present the results of the calculations of Floquet quasienergies using the extended system approach of Sec. IV for the case of the material BC2N.
@@ V.1 Technical details
We have conducted an ab-initio analysis for this material. The first step consisted of a density functional theory calculation using the Quantum Espresso [55] package. We employed a $10\times 10\times 10$ k-mesh both for the self-consistent calculation of the ground state as well as for the subsequent non-self-consistent calculation of the energy eigenvalues and Bloch eigenfunctions. The core-valence interaction was treated by means of scalar relativistic projector-augmented-wave pseudopotentials that had been generated with the Perdew-Burke-Ernzerhof exchange-correlation functional. The pseudopotentials were taken from PS library, generated using the atomic code by A. Dal Corso v.6.3, and the energy cutoff for the plane-wave basis expansion was set at 70 Ry. In the next step, we constructed maximally localized Wannier functions using the Wannier90 package [56]. After discarding the two lowest-lying bands, we considered a set of eight disentangled Wannier functions around the Fermi level to span the “inner” energy range [57] $[-3.33,3.35]\;\text{eV}$. As initial projections, we considered $p_{z}$ atom-centered orbitals for all atoms.
In order to calculate the quasienergy spectrum $\varepsilon_{n}(\bm{k})$, we implement the time discretization method detailed in Appendix C, with matrix representation of the Hamiltonian $\hat{H}(t_{j})$ given by the different methods discussed in Sec. IV.2: the length gauge, Eq. (48), and the velocity gauge Eq. (49) keeping terms up to $M=1$, $M=2$ and $M=3$. We have performed a time complexity analysis of the algorithm in Appendix F, where we have studied the computational cost of the different methods presented in the manuscript.
@@ V.2 Band structure
Figure 3: Band structure of mirror-symmetric BC2N along the $k_{z}=0$ $\Gamma-\text{X}-\text{S}-\text{Y}-\Gamma$ path in the BZ and in the $[-9.0,11.5]\;\text{eV}$ energy range centered around the Fermi energy. The displayed bands have a definite parity with respect to the $\mathcal{M}_{x}$ point group operation, $A_{n}$, which is displayed as a colored over line. The red (blue) overline represents the $+1$ ($-1$) eigenvalue.
@@ V.3.1 General properties and quality of approximations
Figure 4 shows the quasienergy spectrum of the eight bands considered in Fig. 3 as a function of the electric field amplitude. The calculations were performed at k-point $\bm{k}=(1/2,1/4,0)$ (crystal coordinates) lying along the X-S line, and fixed frequency $\hbar\omega=0.5\;\text{eV}$, which corresponds to an almost resonant excitation between the highest occupied molecular orbital (HOMO) and the lowest unoccupied molecular orbital (LUMO). The figure includes three results calculated in the velocity gauge obtained by keeping a different number of expansion terms in Eq. (49); up to $M=1$, $M=2$, and $M=3$. In addition, we have also included results calculated using the length gauge approximation of Eq. (48), employed in other works.
Figure 4 distinguishes three regimes as a function of $E$, which we have separated by vertical lines for the case of the velocity gauge results as a guide to the eye. The first is the noninteracting regime of small $E$, where quasienergy bands are virtually flat. The second is an intermediate electric field regime where the onset of dispersion takes place and is the physically most interesting one. Among other features, this regime shows wide, pronounced avoided crossings that mark multiphoton resonances, which we study in detail in Sec. V.3.2 below. The third regime is the continuum limit of large $E$, which corresponds to $\lambda\gg 1$ [c.f. Eq. (50)] wherein transitions between virtually all energy levels are possible, yielding a featureless continuum spectrum.
@@
V.3.2 Avoided crossings, multiphoton resonances, and generalized parity
@@
where $A_{n}$ is the parity with respect to the $\mathcal{M}_{x}$ operation, as introduced in the discussion of the band structure in Fig. 3. In the case of the HOMO and LUMO states we have $A_{\text{HOMO}}=-A_{\text{LUMO}}=-1$, while the representatives are $m=14$ for the HOMO and $m^{\prime}=15$ for the LUMO state (notice that translation between quasienergy first Brillouin zone is accompanied by a shift in the representative in Fig. 4), hence $P_{\text{HOMO},14}=P_{\text{LUMO},15}$. From the group theory perspective, the quasienergies corresponding to these states cannot cross and therefore undergo an avoided crossing that marks a one-photon resonance since $|m-m^{\prime}|=1$.
@@ VI Summary and Outlook
In this work, we have studied two central quantities of the Floquet theory, namely the effective Floquet Hamiltonian and its eigenvalues, the quasienergies. In the first part of the work, we have analyzed the implications of the electromagnetic gauge freedom in the transformation properties of these quantities. This subject has attracted relatively little attention in the literature so far, and it appears to have been only studied using perturbation theory up to second order in the external field [59, 60]. Here we have explicitly proved the gauge-covariance of the effective Floquet Hamiltonian, which ensures the gauge-invariance of the quasienergy spectrum to all orders. We have provided a numerical example of the latter by discussing the results obtained in the velocity and length gauges for a particle trapped in a one-dimensional box subjected to a periodic drive.
@@ VII Acknowledgments
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the European Research Council (ERC) Grant Agreement No. 946629, Grants No. PID2021-129035NB-I00 and PID2021- 126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”, and from the Basque Government through Grant No. IT1470-22.